supply price = 50+q/2 demand price 150-q/5 find producer and consumer surplus


1
Expert's answer
2021-03-25T13:58:36-0400

Find the equilibrium point


"50+\dfrac{q}{2}=150-\dfrac{q}{5}"

"q=\dfrac{1000}{7}, p=50+\dfrac{500}{7}=\dfrac{850}{7}"

The equilibrium point "(\dfrac{1000}{7}, \dfrac{850}{7})"


The consumer surplus is


"\displaystyle\int_{0}^{1000/7}(150-\dfrac{q}{5})dq-\dfrac{1000}{7}(\dfrac{850}{7})"

"=[150q-\dfrac{q^2}{10}]\begin{matrix}
 1000/7 \\
 0
\end{matrix}-\dfrac{850000}{49}="

"=\dfrac{100000}{49}\approx2040.82"


The producer surplus is


"\dfrac{1000}{7}(\dfrac{850}{7})-\displaystyle\int_{0}^{1000/7}(50+\dfrac{q}{2})dq"

"=\dfrac{850000}{49}-[50q+\dfrac{q^2}{4}]\begin{matrix}
 1000/7 \\
 0
\end{matrix}"

"=\dfrac{250000}{49}\approx5102.04"